Question 4 2 pts Determine whether the vector u is in the column space of the matrix A and whether it is the null space of A.Question 5 1 pts 1 co 2 The vectors u = -3 v= 8 W = -2 is a basis for R3. 3 8 - 7 True False
Answer #1

A- 3 -4 -4 -9 น - 3 3 - 1 .3 4 3 6 1 ***BB HB (1) (2) =), +3C3= -4 -261-62-403= -1 39,-362 = 3 -C, +362+6C3=4 6) From (2) & (a 4x3 matrix , to the null u= does not belongs Since here A is So -] 3 4 4x1 space of A. The hall space of A contan the reaso2 5. .3. w น- د وا ر -2 - ش سه 8 8 -7 3 3 2 1 Now -2 (-56+16) +3 (21+5) 24-24-24) -3 +2 3 -7 8 =-40+81-96 cinearly =-55 +0 si